# Emerging Technologies for 6G Non-Terrestrial-Networks: From Academia to Industrial Applications

### v1

@@ V ORAN-AIDED NTN
TABLE IV: ORAN interface notations
@@
E2
This interface establishes a connection between the near-RT RIC and multiple O-CU-CPs, O-CU-UPs, and O-DUs.
[102]
E1
Interface linking the SMO with O-RAN managed elements
[102]
O1
The interface facilitating communication between management entities (NMS/EMS/MANO) and O-RAN managed elements enables operational and managerial functions. This encompasses FCAPS management (fault, configuration, accounting, performance, and security), software management, and file management
[102]
O2
Interface between the SMO and the O-Cloud
[103]
A1
The interface linking the non-RT RIC and Near-RT RIC facilitates policy-driven guidance of Near-RT operations
[103]
F1-u
F1 User Plane interface
[104]
F1-c
F1 Control Plane interface
[104]
NG-c
The next generation (NG) control plane interface is defined between a NG-RAN node and a 5GC
[105]
NG-u
The NG user plane (NG-U) interface is established between an NG-RAN node and a UPF. It ensures the non-guaranteed delivery of PDU Session/MBS session user plane PDUs between the NG-RAN node and the user plane function (UPF).
[105]
Xn-u
The Xn user plane (Xn-U) interface is established between two NG-RAN nodes. It facilitates the non-guaranteed delivery of user plane PDUs between these two NG-RAN nodes
[106]
Xn-c
Reference point for the control plane protocol interconnecting NG-RAN nodes
[106]
X2-c
Control plane interface for the exchange of application-level configuration data required for seamless interoperability between eNB and en-gNB
[107]
Figure 19: ORAN-aided NTN architecture for the scenario gNB on Earth.
Figure 20: ORAN-aided non-terrestrial network architecture for the scenario gNB on the satellite.
@@
As open networks move the Radio Access Network (RAN) to the cloud, an intriguing alternative is venturing into space. The concept of Non-Terrestrial Networks (NTN) has garnered increasing attention, with numerous experts asserting that this innovative approach holds the potential to extend 6G coverage ubiquitously [109]. Nevertheless, there is currently a scarcity of research in this domain. Consequently, this section delves into the potential utilization of ORAN (Open RAN) for Non-Terrestrial Networks (NTN). We commence by examining the advantages of ORAN, followed by the presentation of a prospective ORAN application architecture tailored for NTN.
V-A Benefits of ORAN for 5G and Beyond
@@ V-B ORAN Architecture
In the description above, we provide an overview of each ORAN function but it would be remiss if we did not mention the interface between them. The details of interfaces in ORAN architecture can be described in Table IV.
V-C ORAN-Aided NTN Architecture for Transparent Satellite, without GNB Onboard
@@
In Fig. 19, we propose a novel scenario while gnB is assumed to be allocated on the Earth. Therefore, it is assumed that the O-DsU is located on Earth and is connected to the NTN gateway (NTNGW) through the next generation (NG) interface [111]. The NTNGW and the UE then connect to the satellite through NG over satellite radio interface (SRI) and NR-Uu interface, respectively [111]. Moreover, the O-CU-CP and O-CU-UP have F1-c and F1-u interfaces with 5GC, respectively.
V-D ORAN-Aided NTN Architecture for Regenerative Satellite, with GNB Onboard
@@
In Fig. 20, we show a novel architecture for the regenerative satellite [112]. To the best of our knowledge, this is the first proposed architecture that shows how ORAN can work with NTN systems [112]. Specifically, the gNB is equipped with a satellite. In this scenario, the NTNGW connects directly to the O-CU-UP through the NG interface. Besides, the NTNGW also connects with non-real-time RIC, and the in-band controls such as E2, F1-c, and F1-u. Then, the NTNGW connects to the satellite through NG over SRI. In this case, the satellite is assumed to be equipped with NTN O-DU. The UE also connects to the satellite through NR-Uu as in the transparent satellite scenario.
@@
V-E ORAN-Aided UAV Trajectory Design Based on Radio Resource Allocation
@@
V-F Flow Diagram for ORAN-Aided UAV Trajectory Design Based on Radio Resource Allocation
@@
VI CHALLENGES, OPEN ISSUES, AND FUTURE RESEARCH DIRECTIONS
VI-A Challenges and Open Issues
VI-A1 Complexity of network slicing
@@
The complexities of constrained optimization problems inherent in network slicing for NTN pose significant challenges. Particularly, these optimization problems are characterized by numerous practical constraints, ranging from bandwidth limitations and latency requirements to security considerations. Moreover, the combination of various platform in different altitudes, e.g., sattellites, UAV, and base station, further aggravates the problem complexity. Consequently, to ensure the effectiveness of network slicing in NTN, optimization techniques need to be carefully designed to take into account the unique characteristics of each platform and the dynamic nature of the network environment.
VI-A2 Challenges of AI-enabled NTN
@@
AI techniques in NTN face a wide range of challenges due to the distinctive features of non-terrestrial communication environments. First, the dynamic nature of NTN, influenced by atmospheric conditions and connectivity shifts, hinders the effectiveness of conventional AI techniques with limited domain adaptation capability. Moreover, constrained bandwidth further complicates the deployment of AI models, limiting the transmission of large datasets, which in turns reduce the effectiveness of AI models’ training. Non-terrestrial platforms also often have limited power and computational capabilities, which complicates the deployment of AI techniques. Additionally, applications demanding high prediction accuracy, such as autonomous driving, introduce challenges in adapting AI models to the inherent propagation delay of NTN. To address these challenges, techniques such as TL can be employed to improve the domain adaption ability of AI models, as well as to address the limited power and computational issues. Moreover, both FL and TL can be utilized to reduce the data transmission demands.
VI-A3 Security
@@
There is a wide range of security issues in NTNs, ranging from data transmission and signal jamming to unauthorized access. Particularly, due to the vast distance of transmission, NTNs are susceptible to interception and eavesdropping. Moreover, the expansive coverage areas of NTN also make them potential targets for unauthorized access. This also brings forth privacy concerns as sensitive data are transmitted over long distances. Aditionally, the reliance on satellite communication introduces challenges related to secure satellite operation, such as satellite commands and cyberattacks. To overcome these problems, promising solutions such as AI-empowered intrusion detection system, blockchain-based authentication, and privacy-preserving homomorphic encryption should be investigated.
VI-B Future Research Directions
VI-B1 Blockchain for NTN
@@
Potential applications of blockchain in NTN encompass diverse areas, aiming to enhance security, interoperability, efficiency, and privacy. For example, by leveraging public/private key and digital signatures mechanisms, blockchain can ensure secure and privacy-preserving identity management in NTN, enabling transparency management and preventing unauthorized access. Moreover, smart contracts can be leveraged to orchestrate network slicing, enabling the automatic execution of agreements among network entities, as well as streamlining the process of allocating and managing resources dynamically. Blockchain can also play a pivotal role in stimulating collaboration and resource-sharing through digital tokens and incentive mechanisms. Particularly, participants can be incentivized to contribute resources, share data, or engage in collaborative efforts through token-based rewards, thereby alleviating resource and data demands in NTN applications.
VI-B2 FL for NTN
@@
In NTN, FL can enable collaborative model training across various non-terrestrial platforms. This collaborative learning paradigm brings two significant advantages. First, data can be used to train the models locally, and thus reducing the risks of exposing sensitive data over the network. Moreover, this also reduces the needs to transmit data, thereby alleviating the transmission demands. Nevertheless, the integration of FL in NTN involves challenges related to varying connectivity, platform mobility, and heterogeneous datasets. Specifically, different non-terrestrial platforms may experience intermittent or fluctuating network connections, affecting the synchronization and coordination required for FL processes. Moreover, the mobility of these platforms introduces complexities in maintaining consistent and reliable communication channels for collaborative model training. Addressing these challenges, therefore, is crucial for unlocking the full potential of FL in NTN and requires more efforts from the academia.
VI-B3 Generative AI for NTN

### v2

@@ 5 ORAN-aided NTN
Table 5: ORAN interface notations
@@
E2 This interface establishes a connection between the near-RT RIC and multiple O-CU-CPs, O-CU-UPs, and O-DUs. [125]
E1 Interface linking the SMO with O-RAN managed elements [125]
O1 The interface facilitating communication between management entities (NMS/EMS/MANO) and O-RAN managed elements enables operational and managerial functions. This encompasses FCAPS management (fault, configuration, accounting, performance, and security), software management, and file management [125]
O2 Interface between the SMO and the O-Cloud [126]
A1 The interface linking the non-RT RIC and Near-RT RIC facilitates policy-driven guidance of Near-RT operations [126]
F1-u F1 User Plane interface [127]
F1-c F1 Control Plane interface [127]
NG-c The next generation (NG) control plane interface is defined between an NG-RAN node and a 5GC [128]
NG-u The NG user plane (NG-U) interface is established between an NG-RAN node and a UPF. It ensures the non-guaranteed delivery of PDU Session/MBS session user plane PDUs between the NG-RAN node and the user plane function (UPF). [128]
Xn-u The Xn user plane (Xn-U) interface is established between two NG-RAN nodes. It facilitates the non-guaranteed delivery of user plane PDUs between these two NG-RAN nodes [129]
Xn-c Reference point for the control plane protocol interconnecting NG-RAN nodes [129]
X2-c Control plane interface for the exchange of application-level configuration data required for seamless interoperability between eNB and en-gNB [130]
Open FH M-Plane Open fronhaul M-Plane interface which is a management interface controlling O-RU [131]
NGAP Control plan protocol of 5G [132]
GTP GTP tunnel to control the UE’s data traffic [132]
NR Uu Satellite radio interface (SRI) between satellite and UE/NTN gateway. [133]
NG over SRI Satellite radio interface between satellite and NTN gateway (NTNGW) in the regenerative satellite scenario. [133]
NG Next generation interface between satellite and gNB/5G core network (5G CN) in the transparent/regenerative satellite scenario. Besides, it is also the interface between gNB and 5G CN in the transparent satellite case. [133]
N6 It is the interface connecting 5G CN to the data network. [133]
Figure 19: ORAN-aided NTN architecture for the transparent NTN and gNB on Earth.
Figure 20: ORAN-aided NTN architecture for the regenerative NTN with gNB onboard.
Figure 21: ORAN-aided non-3GPP NTN architecture.
@@
As open networks move the Radio Access Network (RAN) to the cloud, an intriguing alternative is venturing into space. The concept of NTN has garnered increasing attention, with numerous experts asserting that this innovative approach holds the potential to extend 6G coverage ubiquitously [135]. Nevertheless, there is currently a scarcity of research in this domain. Consequently, this section delves into the potential utilization of ORAN for NTN. In this section, we describe the 3GPP and non-3GPP-based NTN architectures. Our purpose is to provide a comprehensive overview of the NTN architectures from both industrial and academic perspectives. Subsequently, based on these architectures, we will demonstrate how ORAN connects to the NTN system according to each specific one. In particular, this work provides detailed information about the NTN architecture in both 3GPP and non-3GPP environments. Additionally, it offers insights into components within the ORAN system that can connect to satellites, which were previously not clearly addressed. This comprehensive perspective aids researchers in gaining a complete insights into the subject. We commence by examining the advantages of ORAN, followed by the presentation of a general ORAN application architecture. Then, we present each NTN architecture and the corresponding ORAN-based NTN architecture. More specifically, we show two main NTN architectures in 3GPP: transparent and regenerative NTN systems in Release 16 of 3GPP[133]. Then, we show 1 more NTN architecture in non-3GPP scenario. Based on each NTN system, we detail which components in ORAN can connect to the NTN system, which has not been covered before. This can provide an easy-to-understand perspective for academic researchers instead of reading jargon-filled and confusing industry documents.
5.1 Benefits of ORAN for 5G and beyond
@@ 5.2 ORAN architecture
In the description above, we provide an overview of each ORAN function but it would be remiss if we did not mention the interface between them. The details of interfaces in ORAN architecture can be described in Table 5.
5.3 ORAN-aided 3GPP NTN architecture
In this subsection, we describe about 3GPP-based NTN architectures [133]. Then, we outline the specific ORAN components that can interface with it, a perspective not previously presented in detail. In 3GPP Release 16, two primary NTN architectures are listed: transparent and regenerative NTN systems [133]. Consequently, we’ll illustrate these two NTN architectures alongside their respective ORAN counterparts.
5.3.1 ORAN-aided NTN architecture for transparent satellite, without gNB onboard
Fig. 19 a) shows the transparent NTN system whereas the satellite acts as a relay between the NTN ground gateway and the UE [133]. More specifically, the satellite plays the roles of RF amplifier and frequency conversion [133, 137]. Besides, the satellite can replicate the radio interface, transmitting signals from the feeder link (connecting the NTNGW and the satellite) to the service link (linking the satellite and the UE), and vice versa. The interfaces in a transparent NTN system can be described as follows:
  • The UE connects to the satellite through NR Uu interface. Moreover, the NR Uu interface will also be used to connect between satellite and NTNGW and from NTNGW to gNB.
  • The gNB connects to the 5G CN through next generation (NG) interface.
  • N6 interface is used to connect between 5G CN to the data network.
@@ 5.3.1 ORAN-aided NTN architecture for transparent satellite, without gNB onboard
In Fig. 19 b), we show a scenario with transparent NTN while gnB is assumed to be allocated on the Earth [133]. Therefore, it is assumed that the NTN O-DU is located on Earth and is connected to the NTNGW through the next generation (NG) interface [137]. The UE then connects to the satellite through NR-Uu interface [137]. Moreover, the O-CU-CP and O-CU-UP have NGNA and GTP interfaces with 5GC, respectively.
5.3.2 ORAN-aided NTN architecture for regenerative satellite, with gNB onboard
Fig. 20 a) and b) present the regenerative NTN systems in which the satellite is equipped as a base station [133]. In Fig. 20 a), we depict the scenario where only a single satellite connects to the NTNGW. Fig. 20 b) presents a generalized version of Fig. 20 a), incorporating an inter-satellite link (ISL) interface between satellites. The interfaces in regenerative NTN system can be described as follows:
  • The UE connects to the satellite through NR-Uu interface.
  • The Xn interface, operating over the ISL, facilitates connections between satellites. The ISL can utilize either a radio interface (RI) or an optical interface (OI) and may conform to either 3GPP or non-3GPP standards.
  • The regenerative satellite connects to the NTNGW through the NG satellite radio interface (SRI), which serves as a transport link. Moreover, the NTNGW is a transport network layer (TNL) node capable of supporting all transport protocols.
  • N6 interface is used to connect between 5G CN to the data network.
@@ 5.3.2 ORAN-aided NTN architecture for regenerative satellite, with gNB onboard
In Fig. 20 c), we show a novel architecture for the regenerative satellite [138]. To the best of our knowledge, this is the first proposed architecture that shows how ORAN can work with NTN systems [138]. Specifically, the gNB is equipped with a satellite. In this scenario, the NTNGW connects directly to the O-CU-UP and O-CU-UP through the F1-c and F1-u interfaces, respectively. Besides, the NTNGW also connects with non-real-time RIC, and the in-band controls such as E2, F1-c, and F1-u. Then, the NTNGW connects to the satellite through F1/O1/E2 over SRI interface. In this case, the satellite is assumed to be equipped with NTN O-DU. The UE also connects to the satellite through NR-Uu as in the transparent satellite scenario.
5.4 ORAN-aided non-3GPP NTN architecture
In this subsection, we describe one more NTN architecture that is not agreed and listed in 3GPP documents [133]. Then, we show in detail which ORAN components can connect to the NTN system.
In Fig. 21 a), we illustrate a scenario in which the gNB cannot directly connect to the 5G CN [36]. This may occur if the gNB is deployed in an isolated area without an optical link to the 5G CN, or if the connection is disrupted due to a disaster or accident. The interfaces in a transparent NTN system can be described as follows:
  • The UE connects to the gNB through NR Uu interface.
  • The gNB connects to the NTNGW through the NG interface.
  • The satellite connect to the NTNGW and 5G CN through NG over SRI interface.
  • N6 interface is used to connect between 5G CN to the data network.
In In Fig. 21 a), we detail how ORAN components connect to the satellite. In this scenario, the satellite plays as a relay between NTNGW and the 5GC. Therefore, it provides a satellite backhaul link to the NTNGW. Moreover, the satellite connects to the 5GC through NG over SRI interface. As an illustration, let us consider an example to facilitate understanding. Initially, let’s assume there exists an optical link between the 5GC and the gNB. However, in the event of a disaster, this link is destroyed. To expedite the rescue mission in the isolated area, satellites can serve as relays to transmit information from the 5GC to the gNB located in this isolated region.
@@ 5.4 ORAN-aided non-3GPP NTN architecture
5.5 ORAN-aided UAV trajectory design based on radio resource allocation
@@
Recently, there are some works that investigate how ORAN meets UAV communications [139, 140, 141, 142]. In [139], Mitsui et al. tried to solve the fairness problem from ground UEs to UAV by optimizing the UAV’s speed and allocated bandwidth. Particularly, the authors proposed iterative algorithms that run on ORAN architecture to solve the above-mentioned problem. More specifically, they first fixed the allocated bandwidth and optimized the UAV’s flying speed. Then, they used the optimized UAV’s flying speed to optimize the allocated bandwidth. This algorithm was stopped when it reached the saturation point. In [140], Betalo et al. tried to minimize the UAV’s energy consumption by optimizing task scheduling, trajectory design, and resource allocation in wireless sensor networks (WSNs). As the problem is NP-hard, they designed a multi-agent deep reinforcement learning (MADRL) algorithm to solve it, and this algorithm is deployed on the ORAN system. [141] is the first work that investigated the benefits of multi-UAV in ORAN architecture. More specifically, the authors aimed to maximize network utility, i.e., data rate and computing resources, by optimizing the UAVs’ flight paths and the allocated offloading tasks in ORAN architecture. Different to other works that investigated from the theoretical perspective, Lorenzo et al. [142] conducted an experimental study on drone video streaming applications on ORAN architectures. More specifically, the authors designed a control system for UAV positions and transmission directionality to improve the total network performance, i.e., the UE’s uplink throughput. Specifically, they prototyped the proposed solution in a real testbed.
5.6 Flow diagram for ORAN-aided UAV trajectory design based on radio resource allocation
@@
5.7 Summary
In this section, we have discussed multiple approaches for ORAN-aided NTN, i.e., benefits of ORAN for 5G and beyond, ORAN architecture, ORAN-aided 3GPP NTN architecture (ORAN-aided NTN architecture for transparent satellite and ORAN-aided NTN architecture for regenerative satellite), ORAN-aided non-3GPP NTN architecture, ORAN-aided UAV trajectory design based on radio resource allocation. Different to the traditional monolithic RAN architecture, ORAN provides many advantages such as cost reduction, scalability, interoperability between vendors, and disaggregation. It also brings new challenges and opportunities for researchers due to the architecture of ORAN is totally different from the RAN. Some potential research directions of ORAN-aided NTN can be described as follows:
  • Design ORAN-NTN routing path: Designing ORAN-NTN routing paths to efficiently support the heterogeneous requirements of UEs in each area, with different latency and rate requirements and in a large coverage area, has become essential to satisfy the UE’s QoS. This is especially important in emergency/disaster scenarios or when the terrestrial network becomes overloaded.
  • Real time applications: How to design efficient algorithms in an ORAN-NTN system to support real-time (e.g., latency requirement of less than 1 ms), near-real-time (e.g., latency requirement from 10 ms to 1 second), and non-real-time RIC (e.g., latency requirement greater than 1 second) requirements at UEs is a complex issue. Depending on different data rates and latency requirements at UEs, we need to allocate resources and design routing paths efficiently. This problem becomes tricky in large-scale scenarios in practice, e.g., more than 1000 BSs and hundreds of thousands of UEs.
  • AIML for ORAN control: In order to utilize ORAN efficiently on a large scale and automatically, AIML can provide efficient solutions by predicting the incoming traffic of the network. Therefore, it offers good solutions to deal with dynamic control and varying traffic requirements.
  • Security for ORAN-NTN: It is necessary to design, maintain, develop, and test algorithms to improve the security of the ORAN-aided NTN system, especially since the large distance between NTN and ORAN creates high transmission delays, making it vulnerable to attacks
[…]
  • ORAN-based NTN network slicing: How we enable ORAN NTN network slicing to support diverse applications with varying latency and rate requirements. How we design routing paths from NTN through different ORAN components to meet the UE’s QoS requirements with limited resources.
  • NTN-based ORAN digital twin: How to enable ORAN NTN digital twin for real-time, near real-time, and non-real-time in different NTN architectures, such as transparent and regenerative NTN modes.
6 Challenges, Open Issues, and Future Research Directions
6.1 Challenges and Open Issues
6.1.1 Complexity of network slicing
@@
The complexities of constrained optimization problems inherent in network slicing for NTN pose significant challenges. Particularly, these optimization problems are characterized by numerous practical constraints, ranging from bandwidth limitations and latency requirements to security considerations. Moreover, the combination of various platforms in different altitudes, e.g., satellites, UAVs, and base stations, further aggravates the problem complexity. Consequently, to ensure the effectiveness of network slicing in NTN, optimization techniques need to be carefully designed to take into account the unique characteristics of each platform and the dynamic nature of the network environment.
6.1.2 Challenges of AI-enabled NTN
@@
AI techniques in NTN face a wide range of challenges due to the distinctive features of non-terrestrial communication environments. First, the dynamic nature of NTN, influenced by atmospheric conditions and connectivity shifts, hinders the effectiveness of conventional AI techniques with limited domain adaptation capability. Moreover, constrained bandwidth further complicates the deployment of AI models, limiting the transmission of large datasets, which in turn reduces the effectiveness of AI models’ training. Non-terrestrial platforms also often have limited power and computational capabilities, which complicates the deployment of AI techniques. Additionally, applications demanding high prediction accuracy, such as autonomous driving, introduce challenges in adapting AI models to the inherent propagation delay of NTN. To address these challenges, techniques such as Transfer Learning (TL) [143] can be employed to improve the domain adaption ability of AI models, as well as to address the limited power and computational issues. Moreover, both FL and TL can be utilized to reduce the data transmission demands.
6.1.3 Security
@@
There is a wide range of security issues in NTNs, ranging from data transmission and signal jamming to unauthorized access. Particularly, due to the vast distance of transmission, NTNs are susceptible to interception and eavesdropping. Moreover, the expansive coverage areas of NTN also make them potential targets for unauthorized access. This also brings forth privacy concerns as sensitive data are transmitted over long distances. Additionally, the reliance on satellite communication introduces challenges related to secure satellite operation, such as satellite commands and cyberattacks. To overcome these problems, promising solutions such as AI-empowered intrusion detection systems, blockchain-based authentication, and privacy-preserving homomorphic encryption should be investigated.
6.2 Future Research Directions
6.2.1 Blockchain for NTN
Potential applications of blockchain in NTN encompass diverse areas, aiming to enhance security, interoperability, efficiency, and privacy. For example, in [144], a blockchain-based approach is developed for secure data sharing in SAGINs. The proposed approach utilizes session-based authentication and public key cryptography for secure communication between satellites, UAVs, and IoT devices. In this system, a blockchain serves as a decentralized platform to manage authentication processes, leveraging smart contracts to automate and securely execute the authentication protocols between entities.
@@ 6.2.1 Blockchain for NTN
Besides automating the authentication processes, smart contracts can be leveraged to orchestrate network slicing, enabling the automatic execution of agreements among network entities, as well as streamlining the process of allocating and managing resources dynamically. Blockchain can also play a pivotal role in stimulating collaboration and resource-sharing through digital tokens and incentive mechanisms [145]. Particularly, participants can be incentivized to contribute resources, share data, or engage in collaborative efforts through token-based rewards, thereby alleviating resource and data demands in NTN applications.
6.2.2 FL for NTN
@@
In NTN, FL can enable collaborative model training across various non-terrestrial platforms. This collaborative learning paradigm brings two significant advantages. First, data can be used to train the models locally, and thus reducing the risks of exposing sensitive data over the network. Moreover, this also reduces the need to transmit data, thereby alleviating the transmission demands [146]. Nevertheless, the integration of FL in NTN involves challenges related to varying connectivity, platform mobility, and heterogeneous datasets. Specifically, different non-terrestrial platforms may experience intermittent or fluctuating network connections, affecting the synchronization and coordination required for FL processes. Moreover, the mobility of these platforms introduces complexities in maintaining consistent and reliable communication channels for collaborative model training. The transmission of training data for FL, especially over large distances, might also be prone to eavesdropping [147]. Addressing these challenges, therefore, is crucial for unlocking the full potential of FL in NTN and requires more efforts from the academia.
6.2.3 Generative AI for NTN